### v1

# Quantitative Analysis of Performance Drop in DeepSeek Model Quantization

Abstract
Recently, there is a high demand for deploying DeepSeek-R1 and V3 locally, possibly because the official service often suffers from being busy and some organizations have data privacy concerns. While single-machine deployment offers infrastructure simplicity, the models’ 671B FP8 parameter configuration exceeds the practical memory limits of a standard 8-GPU machine. Quantization is a widely used technique that helps reduce model memory consumption. However, it is unclear what the performance of DeepSeek-R1 and V3 will be after being quantized. This technical report presents the first quantitative evaluation of multi-bitwidth quantization across the complete DeepSeek model spectrum. Key findings reveal that 4-bit quantization maintains little performance degradation versus FP8 while enabling single-machine deployment on standard Nvidia GPU devices. We further propose DQ3_K_M, a dynamic 3-bit quantization method that significantly outperforms traditional Q3_K_M variant on various benchmarks, which is also comparable with 4-bit quantization (Q4_K_M) approach in most tasks. Moreover, DQ3_K_M supports single-machine deployment configurations for both NVIDIA H100/A100 and Huawei 910B. Our implementation of DQ3_K_M is released at https://github.com/UnicomAI/DeepSeek-Eval, containing optimized 3-bit quantized variants of both DeepSeek-R1 and DeepSeek-V3.
1 Introduction
DeepSeek-V3 (Liu et al., 2024) and DeepSeek-R1 (Guo et al., 2025) have emerged as state-of-the-art open-source language models since their initial release, achieving top-tier performance across multiple LLM benchmarks  [^1] https://lmarena.ai/?leaderboard. Their combination of exceptional reasoning capabilities and open accessibility has driven widespread adoption in both academic and industrial applications, resulting in a significant demand for on-premises deployment. There are two primary factors that motivate this trend:
Service reliability: The official DeepSeek services frequently experience capacity constraints during peak usage periods, creating operational bottlenecks for production systems.
Data governance: Increasing regulatory requirements and organizational privacy policies necessitate on-premises deployment for sensitive applications in scenarios such as healthcare, finance, and government sectors.
Hosting the full version of DeepSeek-V3 or DeepSeek-R1 on a single machine is appealing due to its simplicity and relatively low cost. However, 671 billion FP8 parameters exceed the device memory available on a typical single machine with 8 GPU/NPU devices (like Nvidia A100/A800/H100/H800/H20 and Huawei Ascend 910B).
Model quantization has emerged as a critical technique for efficient deployment, which helps reduce model memory consumption and enables execution on more affordable hardware configurations. For example, with 4-bit quantization (Q4), the memory cost of DeepSeek-R1’s model weights (not including the KV cache and auxiliary memory for inference) is reduced from 670GB to about 370GB, which could support single-machine deployment for most popular device types. However, it is unclear what the performance of DeepSeek-R1 and V3 will be after being quantized.
In this technical report, we perform a quantitative analysis of the effectiveness of DeepSeek model quantization. Our goal is to answer the following questions.
1. How significant is the performance drop in the quantized DeepSeek models compared to the full-precision versions?
2. Among the full model, the distilled models, and the quantized ones, which version should be deployed for a specific hardware configuration?
To investigate the answers to the above questions, we conducted a quantitative evaluation of quantized DeepSeek series of models that simultaneously examines:
  • Cross-Domain Consistency: Performance impacts across mathematical reasoning (MATH, AIME), code generation (MBPP, LiveCodeBench), and general knowledge (MMLU, C-Eval).
  • Multi-Bitwidth Analysis: 2/3/4/8-bit configurations for quantization.
  • Full-Scale Model Coverage: Comprehensive evaluation of both distilled (32B) and full-parameter (671B) DeepSeek variants.
Through rigorous evaluation, we found that the quantized DeepSeek model retains strong performance, with 4-bit quantization results often comparable to FP8 in many scenarios, demonstrating the high cost-effectiveness of quantized models. We also identified significant potential in dynamic quantization techniques.
Furthermore, by drawing insights from existing quantization techniques, we propose a dynamic 3-bit quantization method (DQ3_K_M) that outperforms the 3-bit quantization implementation in llama.cpp and achieves performance comparable to 4-bit quantization across multiple benchmarks.
Our contributions in this technical report are summarized as follows:
1. We conduct comprehensive evaluation of quantized DeepSeek series models. To the best of our knowledge, this is the first work in the industry to assess quantization effects on full-parameter DeepSeek models. We hope that this work can provide some reference for practitioners who aim to implement DeepSeek models in production environments.
2. We propose a dynamic 3-bit quantization method validated on full-capacity DeepSeek R1 and V3 models, which achievied strong performance. The quantized models (281G) can be conveniently deployed on a single 8 GPU/NPU device (e.g., H100 or 910B). To facilitate community use, we have open-sourced our 3-bit quantized DeepSeek models  [^2] https://github.com/UnicomAI/DeepSeek-Eval.
In the following sections, we first review related works in Section 2, followed by Section 3 where we introduce our proposed dynamic 3-bit quantization method (DQ3_K_M). Our controlled experiments, quantitative results, and practical recommendations for deployment scenarios are presented in Section 4, and finally, we conclude this work in Section 5.
2 Related work
In this section, we provide a brief overview of two techniques for LLM compression: distillation and quantization.
2.1 Distillation
Knowledge distillation (KD) (Hinton et al., 2015; Romero et al., 2014), initially proposed for developing compact yet powerful models through knowledge transfer, has evolved into a fundamental paradigm for model compression (Xu et al., 2024). Traditional KD implementations primarily operate through Logit-level alignment (Hinton et al., 2015) or Intermediate feature matching (Romero et al., 2014).
Recent works (Yang et al., 2024) in LLM distillation demonstrate that supervised fine-tuning (SFT) with teacher-generated outputs presents a viable alternative to conventional KD approaches. Empirical studies (Min et al., 2024; Qin et al., 2024; Huang et al., 2024) have validated that this data-driven distillation paradigm enables parameter-efficient LLMs to attain competitive reasoning performance while maintaining computational tractability.
DeepSeek has assorted to the 800K training data of DeepSeek-R1 to perform SFT on the Qwen and Llama series of models, creating a series of distilled reasoning models (Guo et al., 2025). We select the 32B version (DeepSeek-R1-distill-Qwen-32B) as a representative for evaluation.
2.2 Quantization
Quantization constitutes a fundamental paradigm for model compression, reducing memory footprint by encoding parameters in low-precision representations (Gholami et al., 2022). Contemporary implementations adopt two principal strategies:
Quantization-Aware Training (QAT) QAT (Esser et al., 2019) integrates quantization constraints during full model retraining. While specialized adaptations like LLM-QAT(Liu et al., 2023b) and EdgeQAT (Shen et al., 2024) demonstrate effectiveness for moderate-scale language models, their prohibitive GPU memory demands and extended training cycles render them impractical for large-scale LLMs.
Post-Training Quantization (PTQ) As a computationally efficient alternative, PTQ (Cai et al., 2020) converts pre-trained models to fixed-point representations without revisiting base model training. This approach requires only lightweight parameter calibration (typically ¡0.1% of original training cost) through:
$\min_{\theta}\mathbb{E}_{x\sim\mathcal{D}_{\text{calib}}}\|f_{\text{FP}}(x)-f_{\text{quant}}(\theta,x)\|,$ (1)
where $\mathcal{D}_{\text{calib}}$ denotes the calibration dataset and $\theta$ represents the quantization scales.
PTQ techniques can be further divided into weights-only quantization and weight-activation quantization. Weights-only quantization, such as GPTQ (Frantar et al., 2022) and SpQR(Dettmers et al., 2023), focuses on minimizing precision loss by adjusting weight bit-widths and applying scale transformations to preserve critical weight distributions. Weight-activation quantization (Xiao et al., 2023; Dettmers et al., 2022) compresses both weights and activations, utilizing techniques such as mixed-precision decomposition and channel-wise scaling to achieve an ideal compression rate with less accuracy degradation.
The quantization evaluation results in this report are all based on the weighted-only PTQ paradigm. Although preliminary efforts have been made in quantized LLM evaluation for reasoning tasks, existing studies predominantly focus on single-domain evaluations (e.g., either mathematical reasoning (Li et al., 2025) or code generation (Giagnorio et al., 2025; Nyamsuren, 2024)). Besides, current analysis (Liu et al., 2025) about DeepSeek quantization are mainly restricted to parameter-constrained distilled variants of the DeepSeek family (less than 32B).
Unlike existing studies, our work introduced in this technical report presents the first systematic study of multi-bitwidth quantization effects across the complete DeepSeek model spectrum, including the full-parameter R1 and V3 variants (671B).
3 Methodology
While existing quantization implementations for DeepSeek models demonstrate preliminary success, we posit that dynamic bit-width allocation [^3] https://unsloth.ai/blog/deepseekr1-dynamic based on layer importance warrants systematic exploration. Building upon the Q3 quantization baseline, we introduce adaptive precision selection guided by architectural insights. Our dynamic quantization strategy prioritizes applying higher-precision quantization to modules with fewer parameters where possible. Therefore, building upon the standard Q3 quantization provided in llama.cpp, we implement hybrid precision by applying q6_k or q4_k quantization to some selected modules. Furthermore, motivated by (Yu et al., 2024)’s discovery of “super weights” in LLMs - particularly concentrated in the mlp.down_proj layers - we observe that applying overly aggressive quantization strategies to these critical components leads to significant model performance degradation. Therefore, we implement:
  • q6_k quantization for the first two ffn_down_exps layers
  • q3_k for subsequent layers with q4_k inserted every fourth layer
This configuration achieves parameter distribution: 75.9% q3_k, 20.7% q4_k, and 3.4% q6_k within ffn_down_exps module.
The resultant DQ3_K_M variant demonstrates superior memory efficiency compared to some conventional approaches (Table 1). Our implementation achieves smaller model footprint with reduced GPU memory consumption and more effective average bit-width against llama.cpp’s standard Q3_K_M.
Quantitative performance comparisons across reasoning and generation tasks are detailed in Section 4. Complete implementation specifics of DQ3_K_M, including per-module quantization schemes, are provided in Appendix A.1.
Metric Q4_K_M (llama.cpp) Q3_K_M (llama.cpp) DQ3_K_M (ours) Q2_K_L (llama.cpp) UD-Q2_K_XL (Unsloth)
Model Size 377G 298G 281G 228G 212G
Avg Quants 4.82 3.81 3.59 2.91 2.70
MU (total) 568GB 487GB 469GB 415GB 398GB
MU (per GPU) 71GB 61GB 59GB 52GB 50GB
Table 1: Comparison of resource consumption between our proposed DQ3_K_M and various quantization approaches provided by llama.cpp and Unsloth, using DeepSeek R1(671B) as an example. The memory usage is reported based on the maximum context length of 32K tokens. MU denotes Memory Usage.
4 Experiments
4.1 Benchmarks
We conducted experiments across two categories of benchmarks  [^4] Please refer to Appendix A.2 for statistics of these benchmarks.: domain-specific reasoning tasks and general capability assessments. Our reasoning benchmark suite comprises nine components :
MATH 500 (Lightman et al., 2023): A curated subset of 500 competition-level mathematics problems from the MATH dataset Hendrycks et al. (2021);
AIME 2024  [^5] https://maa.org/math-competitions/american-invitational-mathematics-examination-aime: It features problems from the American Invitational Mathematics Examination 2024 which are specifically designed to challenge the top high school students;
GPQA (Rein et al., 2024): A Q&A benchmark containing 198 multiple-choice questions spanning physics, biology, and chemistry;
LiveCodeBench(Jain et al., 2024): Temporal programming challenges collected from competitive coding platforms (AtCoder/LeetCode), maintaining temporal consistency (2024-08 to 2025-01) with DeepSeek-R1’s evaluation protocol (Guo et al., 2025);
MBPP(Austin et al., 2021): MBPP (Mostly Basic Python Programming) is a benchmark for assessing LLM’s ability to generate code for independent Python functions. It consists of 974 entry-level programming problems.
MBPP+ (Liu et al., 2023a): An enhanced variant of MBPP featuring expanded test cases and refined solution specifications.
For general capability evaluation, we adopt three established benchmarks:
MMLU (Hendrycks et al., 2020) (Massive Multitask Language Understanding): widely used benchmark for LLM evaluation contains diverse questions across 57 academic subjects.
CMMLU (Li et al., 2023) (Chinese Massive Multitask Language Understanding): 11582 Chinese questions spanning STEM and humanities.
C-Eval (Huang et al., 2023): 12342 challenging Chinese exam-style questions.
DeepSeek-R1 FP8 (Reported) FP8 (Official API) Q4_K_M (llama.cpp) Q3_K_M (llama.cpp) UD-Q2_K_XL (Unsloth) DQ3_K_M (Ours)
AIME 2024 79.8 77.53 ($\pm$2.97) 75.43 ($\pm$3.07) 72.50 ($\pm$6.11) 75.83 ($\pm$5.83) 75.41 ($\pm$4.69)
MATH 500 97.3 95.45 ($\pm$0.82) 95.55 ($\pm$0.44) 94.15 ($\pm$0.68) 95.25 ($\pm$0.44) 95.35 ($\pm$0.50)
GPQA 71.5 69.58 ($\pm$1.65) 69.95 ($\pm$1.85) 65.80 ($\pm$2.30) 68.93 ($\pm$1.55) 68.95 ($\pm$0.65)
MBPP - 92.60 ($\pm$0.80) 91.60 ($\pm$2.00) 90.43 ($\pm$0.88) 92.93 ($\pm$0.24) 92.80 ($\pm$0.70)
MBPP+ - 78.35 ($\pm$1.06) 76.70 ($\pm$1.85) 76.75 ($\pm$0.88) 78.33 ($\pm$0.91) 78.60 ($\pm$1.01)
LiveCodeBench 65.9 64.16 ($\pm$1.51) 62.41 ($\pm$2.27) 61.95 ($\pm$1.66) 61.40 ($\pm$1.59) 63.15 ($\pm$1.06)
MMLU 90.8 90.99 90.14 89.87 89.72 91.03
CMMLU - 90.37 90.42 89.85 89.61 90.17
C-Eval 91.8 92.20 92.10 91.60 91.70 91.80
Average - 83.48 82.70 81.44 82.63 83.03
Weighted avg. - 85.82 85.24 84.28 85.02 85.53
Accuracy drop - - 0.68% 1.80% 0.94% 0.34%
Table 2: Main results of DeepSeek-R1 on various benchmarks. Accuracy drop refers to the relative percentage decrease in average score against the results from FP8 (Official API).
DeepSeek-V3 FP8 (Reported) FP8 (Tencent API) Q4_K_M (llama.cpp) Q3_K_M (llama.cpp) Q2_K_L (llama.cpp) DQ3_K_M (Ours)
AIME 2024 39.2 38.34 ($\pm$2.52) 41.66 ($\pm$4.72) 38.73 ($\pm$4.70) 15.41 ($\pm$3.55) 39.16 ($\pm$4.97)
MATH 500 90.2 89.85 ($\pm$0.30) 90.55 ($\pm$0.44) 89.05 ($\pm$1.27) 77.30 ($\pm$0.66) 89.65 ($\pm$0.98)
GPQA 59.1 52.23 ($\pm$3.44) 51.95 ($\pm$2.64) 52.13 ($\pm$1.25) 43.65 ($\pm$1.32) 52.38 ($\pm$1.31)
MBPP - 87.75 ($\pm$0.61) 87.18 ($\pm$0.70) 88.55 ($\pm$0.90) 81.10 ($\pm$1.55) 89.38 ($\pm$0.35)
MBPP+ - 73.35 ($\pm$1.21) 72.90 ($\pm$0.66) 73.08 ($\pm$1.31) 67.83 ($\pm$1.09) 74.78 ($\pm$0.56)
LiveCodeBench 36.2 36.21 ($\pm$0.47) 37.40 ($\pm$1.32) 36.21 ($\pm$2.03) 29.14 ($\pm$0.92) 36.76 ($\pm$0.67)
MMLU 88.5 88.06 88.09 87.31 84.25 87.87
CMMLU - 81.57 82.68 80.69 77.32 81.07
C-Eval 86.5 83.10 82.90 82.60 77.60 83.40
Average - 70.05 70.59 69.82 61.51 70.47
Weighted avg. - 75.45 75.79 75.06 68.73 75.73
Accuracy drop - - 0 0.52% 8.91% 0
Table 3: Quantization results of DeepSeek-V3 on various benchmarks.
DeepSeek-R1 distill-Qwen-32B BF16 (Reported) BF16 (Local Evaluation) Q8_0 (llama.cpp) Q4_K_M (llama.cpp) Q3_K_M (llama.cpp)
AIME 2024 72.6 69.59 ($\pm$2.75) 71.68 ($\pm$4.71) 70.40 ($\pm$7.66) 71.24 ($\pm$6.66)
MATH 500 94.3 93.65 ($\pm$0.41) 93.10 ($\pm$0.42) 93.90 ($\pm$0.53) 93.50 (0.38)
GPQA 62.1 61.85 ($\pm$2.18) 58.85 ($\pm$2.75) 62.00 ($\pm$4.54) 60.20 ($\pm$1.95)
LiveCodeBench 57.2 57.08 ($\pm$1.01) 57.59 ($\pm$1.17) 56.85 ($\pm$2.87) 55.20 ($\pm$1.74)
MBPP - 89.35 ($\pm$0.42) 89.35 ($\pm$0.73) 89.73 ($\pm$1.20) 88.93 ($\pm$0.64)
MBPP+ - 75.43 ($\pm$0.91) 75.45 ($\pm$1.18) 75.53 ($\pm$1.04) 75.38 ($\pm$1.30)
MMLU - 82.15 82.15 82.37 82.17
CMMLU - 83.91 83.97 83.57 83.34
C-Eval - 87.0 86.7 86.8 86.2
Average - 77.78 77.65 77.91 77.35
Weighted avg. - 79.94 79.71 79.97 79.40
Accuracy drop - - 0.29% 0 0.68%
Table 4: Results of DeepSeek-R1-distill-Qwen-32B on various benchmarks. Accuracy drop denotes the relative decrease in average score against the results from BF16.
4.2 Experimental Setting
We evaluate the performance of quantized models from three original models (DeepSeek-V3, DeepSeek-R1 and DeepSeek-R1-distill-Qwen-32B) across multiple bit-width configurations on the aforementioned benchmarks. Our post-training quantization (PTQ) implementation leverages two established frameworks:
  1. 1.
    llama.cpp [^6] https://github.com/ggml-org/llama.cpp for 4-bit (Q4_K_M), 3-bit (Q3_K_M), 2-bit (Q2_K), and 8-bit (Q8_0) configurations
  2. 2.
    Unsloth [^7] https://unsloth.ai/blog/deepseekr1-dynamic for specialized dynamic 2-bit quantization (Q2_K_XL)
Quantization Setting
The quantization configurations shared by all models include:
  • 4-bit: Q4_K_M (llama.cpp)
  • 3-bit: Q3_K_M (llama.cpp)
Model-specific quantization implementations:
  • DeepSeek-V3 2-bit: Standard Q2_K (llama.cpp)
  • DeepSeek-R1 2-bit: Large-scale UD-Q2_K_XL (unsloth)
  • DeepSeek-distill-Qwen-32B 8-bit: Q8_0 (llama.cpp)
For DeepSeek-R1 and DeepSeek-V3, we also conduct additional performance evaluations of our proposed Q3 quantization implementation (DQ3_K_M).
Decoding Configuration
All quantized models were configured with a maximum generation length fixed at 32,768 tokens. We used a temperature of 0.6 and a top-p value of 0.95. We implemented differentiated decoding strategies across benchmark categories:
  1. 1.
    For small benchmarks (MATH 500, GPQA, LiveCodeBench, etc.), we employ rigorous statistical sampling: generating 4 independent responses per query and compute mean scores across samples to mitigate variance. Since AIME 2024 only contains 30 questions, we sampled 8 responses for each question.
  2. 2.
    For large benchmarks (MMLU, CMMLU, and C-Eval), we adopt a single inference pass per question, as we observe relatively stable results on these benchmarks.
4.3 Main Results
The evaluation results for DeepSeek-R1, DeepSeek-V3 and DeepSeek-R1-distill-Qwen-32B are shown in Table 2, 3, and 4, respectively. We present the official evaluation results reported in (Guo et al., 2025), official deepSeek API [^8] https://api-docs.deepseek.com/ invocation outcomes, and performance metrics of different quantized model variants. For multi-sampling results, we report mean values with corresponding standard deviations (in parentheses). Notably, due to the official DeepSeek V3 API update on March 24, 2025, we substituted it with Tencent’s DeepSeek V3 API  [^9] https://cloud.tencent.com/document/product/1772/115963 in Table 2 to ensure comparability.
DeepSeek-R1
Table 2 demonstrates the impact of various quantization methods on DeepSeek-R1’s performance across multiple benchmarks. While the official FP8 demonstrates superior overall performance, Q4_K_M quantization methods exhibit surprisingly competitive results. Across most reasoning benchmarks, Q4_K_M shows no significant performance degradation, with its metrics on MATH 500 and GPQA even marginally outperforming the official FP8 API implementation. The performance on general capability benchmarks remains stable across all variants, suggesting core semantic representations withstand quantization.
The proposed DQ3_K_M approach attaining an average score of 83.03 that surpasses standard 3-bit implementations and closely matches the performance of Q4_K_M. These results, in conjunction with Table 1, demonstrate that DQ3_K_M achieves superior efficiency without compromising capability. The performance stability of DQ3_K_M also proves particularly noteworthy, achieving lowest standard deviation in scientific QA (0.65 on GPQA vs Q3_K_M’s 2.30) and coding benchmarks (1.06 on LiveCodeBench vs Q4_K_M’s 2.27). We also noticed that dynamic quantized 2-bit model (UD-Q2_K_XL) outperforms standard 3-bit quantization (Q3_K_M) across multiple benchmarks, further validates the fundamental benefits of dynamic quantization.
DeepSeek-V3
As evidenced in Table 3, the DeepSeek-V3 model exhibits similar quantization characteristics to DeepSeek-R1 in general. Through comparison of standard quantization variants from llama.cpp, we could find that Q3_K_M performs slightly worse than Q4_K_M (with weighted average of 75.06 vs. 75.79). Approach Q2_K_M exhibits severe performance degradation in all evaluated benchmarks (61.51 vs. 70.05 on average when compared with FP8). This empirically validates the inevitable accuracy-compression trade-off in LLM quantization, where aggressive bit-width reduction fundamentally disrupts model capabilities. With a weighted average benchmark score of 75.73, our proposed dynamic 3-bit quantization (DQ3_K_M) performs similarly with Q4_K_M (75.79) and FP8 (75.45). These results again demonstrate the effectiveness of our newly proposed quantization approach.
DeepSeek-R1-distill-Qwen-32B
As reported in Table 4, our systematic evaluation of DeepSeek-R1-distill-Qwen-32B reveals that 4-bit quantization (Q4_K_M) achieves optimal performance preservation, maintaining the performance of the original BF16 format across diverse benchmarks while reducing memory requirements significantly. This configuration demonstrates particular robustness in mathematical reasoning (MATH: 93.90 vs 93.65 local BF16) and scientific QA (GPQA: 62.00 vs 61.85), despite exhibiting higher standard deviation in complex tasks ($\sigma=7.66$ for AIME 2024). For code generation tasks, MBPP and MBPP+ show remarkable quantization resilience performance variation across bit-widths. However, LiveCodeBench shows sensitivity to aggressive quantization (Q3_K_M: 55.20 vs Q8_0: 57.08). This may be due to the relatively high difficulty of LiveCodeBench. In addition, consistent performance preservation ($\varDelta<0.8\%$) across MMLU/CMMLU/C-Eval shows exceptional robustness of different bit-widths, which demonstrates that quantization preserves the general language understanding capabilities of the distillation model.
Metric Q4_K_M (llama.cpp) Q3_K_M (llama.cpp) DQ3_K_M (Ours) Q2_K_L (llama.cpp) UD-Q2_K_XL (Unsloth)
Avg. Score (V3) 75.79 75.06 75.73 68.73 -
Avg. Score (R1) 85.24 84.28 85.53 - 85.02
MU (total) 568GB 487GB 469GB 415GB 398GB
MU (per GPU) 71GB 61GB 59GB 52GB 50GB
Table 5: Comparison among various quantization approaches in terms of accuracy and memory usage. Memory usage (MU) is reported based on the maximum context length of 32K tokens.
4.4 Recommendations for Different Devices
Based on the statistical analysis in Table 5, we conclude that for full-parameter R1 and V3 models, 4-bit quantization (Q4_K_M) and our DQ3_K_M achieve optimal cost-performance ratio under NVIDIA-based single-machine deployments (e.g., 80GB VRAM per A100/A800/H100/H800/H20 GPU). However, Q4 typically exceeds the VRAM constraints of Huawei Ascend 910B single-node configurations (64GB per NPU), whereas DQ3_K_M satisfies both NVIDIA H100 and Ascend 910B configuration. Compared to other quantization variants, our dynamic 3-bit quantization DQ3_K_M achieves a favorable performance-resource trade-off.
5 Conclusion
This work presents the first systematic evaluation of multi-bitwidth quantization for various deepseek models including 671B-scale, employing a comprehensive analysis across multi-domain benchmarks. Our findings demonstrate that standard 4-bit quantization (Q4) exhibits minimal performance degradation versus FP8 while significantly reducing memory requirements. We further introduce DQ3_K_M, a dynamic Q3 quantization method with higher memory compression ratio that surpasses the current state-of-the-art Q3_K_M implementation in llama.cpp, achieves 1.48% and 0.89% improvement on average on R1 and V3, respectively. This study establishes that careful quantization design can retain the vast majority of the original model’s capabilities with only tiny performance loss while enabling cost-effective deployment on single a single machine with 8 GPU devices. We will explore more efficient quantization techniques in the future.
Appendix A Additional details
A.1 Quantization Implementation Details
Weight-Matrix Q4_K_M Q3_K_M DQ3_K_M (ours) Q2_K_L Q2_K_XL
output q6_k q6_k q6_k q6_k q6_k
token_embd q4_k q3_k q4_k q4_k q4_k
attn_kv_a_mqa q4_k q3_k q6_k q6_k q6_k
attn_kv_b q4_k q3_k q6_k q2_k q6_k
attn_output q4_k q4_k q4_k q3_k q4_k
attn_q_a q4_k q3_k q4_k q2_k q4_k
attn_q_b q4_k q3_k q4_k q2_k q4_k
ffn_down q6_k q5_k q6_k q3_k q6_k
ffn_gate q4_k q3_k q4_k q2_k q4_k
ffn_up q4_k q3_k q4_k q2_k q4_k
ffn_down_exps q4_k(53.4%) q6_k(46.6%) q4_k q3_k(75.9%) q4_k(20.7%) q6_k(3.40%) q3_k q2_k(94.8%) q3_k(5.20%)
ffn_down_shexp q4_k(53.4%) q6_k(46.6%) q4_k q6_k q3_k q6_k
ffn_gate_exp q4_k q3_k q3_k q2_k q2_k
ffn_gate_shexp q4_k q3_k q4_k q2_k q4_k
ffn_up_exps q4_k q3_k q3_k q2_k q2_k
ffn_up_shexp q4_k q3_k q4_k q2_k q4_k
Table 6: Quantization implementation details of different methods
A.2 Benchmark Statistics
Benchmark Question Count Weight
MATH 500 500 0.2
AIME 2024 30 0.5
GPQA 198 0.5
MBPP 378 0.5
MBPP+ 378 0.5
LiveCodeBench 272 0.5
MMLU 14042 1
CMMLU 11582 1
C-Eval 12342 1
Table 7: The statistics of benchmarks for evaluation (The weight is used for calculating weighted average scores in experiments.)